OpenStudy (anonymous):

what is 8 2/5 divided by 3

OpenStudy (solomonzelman):

\[\huge\color{red}{8\frac{2}{5}}~~~~->~~~~\huge\color{red}{\frac{(8 \times 5)+2}{5}}~~~~->\]\[\huge\color{red}{\frac{42}{5}}\]new lets divide this by 3.

OpenStudy (solomonzelman):

\[\huge\color{red}{\frac{42}{5} \div 3~~~~->~~~~\huge\color{red}{\frac{42}{5} \times \frac{1}{3}}}\]\[~~~->~~~~~\huge\color{red}{\frac{42}{5 \times 3}}~~~~->~~~~\huge\color{red}{\frac{42}{15}}\]

OpenStudy (solomonzelman):

\[~~~->~~~~\huge\color{red}{\frac{14 \times 3}{5 \times 3}=} \huge\color{blue}{?}\]

OpenStudy (solomonzelman):

14/5

OpenStudy (solomonzelman):

\[\huge\color{green}{\frac{14}{5}}~~~~~~or~~~~~~\huge\color{green}{2~\frac{4}{5}}\]

OpenStudy (anonymous):

another question if that's ok with you

OpenStudy (anonymous):

whats 7 5/6 divided by 1 1/5

OpenStudy (solomonzelman):

\[\huge\color{red}{7\frac{5}{6}}~~->~~\huge\color{red}{\frac{(7 \times 6)+5}{6}}~~->~~\huge\color{red}{\frac{47}{6}}\]

OpenStudy (solomonzelman):

divide by \[\huge\color{red}{1\frac{1}{5}}~~->~~\huge\color{red}{\frac{(1 \times 5)+1}{5}}~~->~~\huge\color{red}{\frac{6}{5}}\]

OpenStudy (solomonzelman):

Formula. \[\huge\color{red}{\div \frac{a}{b}}~~~->~~~\huge\color{red}{\times \frac{b}{a}}\]

OpenStudy (solomonzelman):

So \[\huge\color{red}{\frac{47}{6}} \div \huge\color{red}{\frac{47}{6}}~~->~~\huge\color{red}{\frac{47}{6}} \times \huge\color{red}{\frac{5}{6}}~~->~~\huge\color{red}{\frac{235}{12}}\] reduce,